OpenStudy (anonymous):

Solve. x^2 – x = 20. Answer A. x = –10 and x = 2 B. x = –5 and x = 4 C. x = 5 and x = –4 D. x = 10 and x = –2 Solve. x^2 – x = 20. Answer A. x = –10 and x = 2 B. x = –5 and x = 4 C. x = 5 and x = –4 D. x = 10 and x = –2 @Mathematics

OpenStudy (hoblos):

C

OpenStudy (anonymous):

Thank you, love (:

OpenStudy (anonymous):

\[x^2-x-20=0\] \[(x-5)(x+4)=0\] \[x-4=0\implies x=4\] \[x+5=0\implies x=-5\]

OpenStudy (hoblos):

:D

OpenStudy (anonymous):

in case you actually have to do this on a test, good to have some method right?

OpenStudy (anonymous):

Of course. Lol Thank you :)

myininaya (myininaya):

satellite are you okay?

myininaya (myininaya):

you switched your signs

myininaya (myininaya):

x+4=0 => x=-4 x-5=0 => x=5